# What Changed and What Could Have Changed? State-Change Counterfactuals for Procedure-Aware Video Representation Learning

### v1

@@ 4.1 Implementation Details
Pretraining. We pretrain our model on EgoClip [39], a dataset derived from Ego4D [21], with clean clip-level narrations and the video-level summaries. We follow [39, 2] to use the same training split that has 3.8M clip narrations and 120K long-term video summaries for pertaining. We pretrain our model on 8 NVIDIA L40S GPUs, with batch size 18 on each GPU for 7 epochs. We follow HierVL [2] and train models on clip- and video-level alignments alternatively. Specifically, we perform video-level alignment as in Eq. (5) after every 5 mini-batches of clip-level alignment as in Eq. (4). The frame-level loss is computed during the clip-level training iteration. We set the learning rate and weight decay to 1e-5 and 1e-4, respectively. Note that each short clip is sub-sampled to 4 frames during training [39, 2].
@@
Text Generation. To generate all state-change descriptions and state-change counterfactuals, we use Llama 3-2 8B [15] for efficiency. We incorporate off-the-shelf text-encoder FLAVA [66] to pre-extract features of the annotated narrations and summaries and our generated text descriptions.
@@
Temporal Action Segmentation. We evaluate representations on temporal action segmentation, where models output the frame-wise action label for untrimmed, long videos, requiring the ability to reason about procedures and recognize the start and end of state transformations. We evaluate various representations by adopting the popular temporal action segmentation model ASFormer [90] that takes video features as input. We follow prior work [36] for the same training strategy and evaluation protocol and metrics including F1 scores, Edit distance, and frame-wise Accuracy, on the cooking procedure datasets, Georgia Tech Egocentric Activities (GTEA) [18] and EgoPER [34] Dataset. GTEA consists of 28 egocentric instructional videos capturing daily kitchen activities at 15 frames per second. EgoPER contains 368 videos spanning 28 hours of cooking scenarios.
@@
Error Detection. We also evaluate representations on unsupervised error detection on the EgoPER dataset [34] where models are trained with error-free videos and output labels for each frame on whether it is an error during testing time. The error categories include step omission, step addition, step modification, step slip, and step correction. We follow [34] and use its EgoPED model [34], and equip it with different video representations. Specifically, the training process of EgoPED consists of two stages: action segmentation and prototype learning. In stage 1, the temporal action segmentation model, ASFormer, is trained to predict frame-wise action labels on the error-free set by taking pre-extracted video features as input. In stage 2, the model learns prototype features for each class of action step, e.g., “Take bowl from microwave,”, with contrastive step prototype learning [51, 34]. For each class of activity, multiple prototypes are extracted for error detection at inference. Note that each activity is trained independently, such as Make Coffee. During the inference, EgoPED calculates the similarity score of observed frame features and any learned prototypes. A threshold is further used to determine whether the observed test frame is erroneous.
Baselines. We evaluate the popular I3D and CLIP feature [7, 53] commonly used in long-form video tasks and procedure-aware representations with publicly available pretrained model weights, including MIL-NCE [47], Bridge-prompt [36], and PVRL [98]. In addition, we evaluate the VLM HierVL [2] with only its video encoder.
@@
Table 2: Error detection results on the EgoPER dataset. The underline denotes the second best. HTM denotes the HowTo100M dataset [46]. Text denotes the VLM model with a trainable text encoder.
@@
HF2-VAD [43] - 34.5 62.6 25.4 62.3 29.1 52.7 10.0 59.6 36.6 62.1 27.1 59.9
S3R [82] Kinetics+I3D 52.6 51.8 47.8 61.6 50.5 52.4 16.3 51.0 47.8 57.9 43.0 54.9
I3D [7] Kinetics 62.7 65.6 51.4 65.1 59.6 55.0 55.3 58.3 56.0 66.0 57.0 62.0
CLIP [53] WIT [71]+Text 77.6 67.2 69.6 67.5 66.9 59.5 68.5 69.0 75.6 57.7 71.6 64.2
MIL-NCE [47] HTM 77.3 59.8 69.8 61.5 65.7 53.0 68.0 67.9 69.8 61.5 70.1 60.7
PVRL [98] HTM 75.7 70.0 71.2 53.5 65.5 65.5 67.5 65.4 76.4 65.2 71.3 63.9
HierVL [2] Ego4D+Text 77.9 70.2 70.8 66.4 65.2 58.1 67.4 69.8 75.1 66.4 71.3 66.2
@@ 4.2 Temporal Action Segmentation Results
Table 1 presents results on temporal action segmentation. The first line of the table shows Br-prompt [36] which is pre-trained on the target dataset (GTEA) and thus serves as an upper bound on performance. In our results, we found that MIL-NCE performs significantly worse than others, especially on GTEA. We assume this is because MIL-NCE is trained with low-quality automatic-speech-recognition (ASR) text and imprecise alignment between video and ASR sentences, unlike PVRL [2] which leverages pseudo-labels generated by CLIP [53]. We further compare our model with the VLM HierVL. Even though HierVL is trained with both video and text encoders, which require considerably more computational resources, our model outperforms HierVL in most metrics on both datasets. Furthermore, the performance gap against all non-VLM models is larger. This highlights the effectiveness of our proposed state-change descriptions and their counterfactuals on long-term procedure understanding.
@@ 4.3 Error Detection Results
Table 2 presents results of the error detection benchmark on EgoPER [34]. We first include results reported in EgoPER as a reference in the top group (rows 1-3) instead of a direct comparison. In the remaining groups (rows 4-9), we report the performance of representation learning approaches, tested using the same downstream architecture [90]. We observe that all procedure-aware representations outperform general visual representations, such as I3D and CLIP, highlighting the importance of procedure awareness in the context of error detection. Furthermore, our proposed method surpasses the state-of-the-art in EDA [34], even outperforming the VLM HierVL.
Moreover, we achieve competitive performance on the AUC metric. Note that although we follow prior work [34] to include frame-wise metric AUC, this metric may not be effective in measuring the performance and the procedure awareness of representations. As discussed in EgoPER [34], the heuristic random method can achieve competitive results in AUC but fail to localize erroneous segments as shown in the major metric EDA. This is also true when comparing I3D and earlier methods, such as HF2-VAD [43] and S3R [82]. We hope our extensive experimental results can inspire future work to study more effective evaluation metrics for error detection. In summary, our representation learning method achieves state-of-the-art performance in EDA and competitive performance in AUC, highlighting its strong effectiveness in procedure awareness.
Table 3: Ablation of the proposed state changes and counterfactuals (CF): Before & After states; State-change Counterfactuals (SC-CF); Missing key (K-CF) and Misordered (M-CF) on GTEA for temporal action segmentation and EgoPED for error detection. Note that we calculate EDA by averaging all categories of activity.
Clip-level Video-level GTEA EgoPER
ID Before & After SC-CF K-CF M-CF F1@50 Edit Acc EDA
1 77.2 85.4 77.9 71.9
2 78.5 87.0 78.4 70.5
3 76.6 84.2 77.9 72.5
4 76.7 83.9 78.2 71.7
5 78.4 85.6 78.7 69.0
Ours 81.6 86.8 80.0 72.7
Figure 3: Qualitative results of Error Detection on EgoPER. GT denotes ground truth. Green/Red segmentation and text denote the normal and error labels, respectively. The presented erroneous procedure “Make tea” consists of [Not checking water temperature in the kettle,” Hold the kettle,“Pour water immediately,” “Stir with the knife”, ….].
Figure 4: Qualitative results of Temporal Action Segmentation on EgoPER. The example presents the procedure “Make Pinwheel”. Distinct colored segments are different action step classes.
@@ 4.4 Ablation Study
In Table 3, we study the effectiveness of state changes (Before and After, state-change counterfactuals (SC-CF), and video-level state-change counterfactuals (Missing-step CF and Misordered-step CF) by ablating each of them and evaluating on temporal action segmentation with GTEA and error detection with EgoPER.
Temporal Action Segmentation. One interesting insight on GTEA is that using only clip-level state changes (ID 2) achieves competitive performance on F1 and Edit metrics, which outperforms most other ablated models. Also, using only one type of video-level counterfactual (ID 4 and 5) may mislead the model to overfit a certain type of counterfactual, resulting in lower performance. By considering all the proposed states and counterfactuals together, we show that they are complementary to each other, thus achieving the best performance.
Error Detection. We analyze the influence of the proposed state changes and different types of counterfactuals on error detection. The results on EgoPER show that our proposed state changes and their counterfactuals can improve performance by combining them all together, verifying the effectiveness of recognizing erroneous labels. In addition, we observe interesting results showing that with only clip-level state changes and their counterfactuals (ID 3), the model achieves competitive results in EDA, which highlights its ability to comprehend short-term action-induced transformation and accurately localize erroneous action in frames. We hope our insightful findings can encourage future work to study the strategy for balancing short-term and long-term state-change with counterfactuals for error detection.
4.5 Qualitative Results
Figure 3 presents qualitative results on error detection in EgoPER, showing that our learned representations identify erroneous activities with greater fidelity. PVRL and HierVL produce several false positives, leading to over-segmentation, whereas our method better aligns with the ground-truth segments in both temporal location and count. Figure 4 shows qualitative results on Temporal Action Segmentation, where PVRL and HierVL misclassify large temporal segments (dark and light orange, respectively). In contrast, our model more accurately distinguishes and classifies actions, reflecting an improved understanding of procedural activities and aligning with our quantitative results.
@@ 5 Conclusions
Appendix A Text Description Generation
In this section, we present the text-generation process used in this work. We generate clip-level state-change descriptions, i.e., Before, After and State-change counterfactual; and video-level state-change counterfactuals, i.e., Missing-key-Step Counterfactuals and Misordered-Step Counterfactuals. In this work, we use Llama 3.1 [15], the latest version of Llama at the time of implementation. To generate the text descriptions for the significantly large dataset Ego4D [21], we select Llamma3.1 8B for efficiency.
A.1 Prompt Design
We feed the clip narrations and video summaries to Llamma to generate the corresponding states and counterfactuals. Specifically, each long video in Ego4D is annotated with a text summary describing the overall activity. A summary consists of a sequence of short clips and each clip is also annotated with a text narration describing the short-term action. Below, we present the generation of clip- and video-level texts separately:
Clip Level
Before, After and State-change counterfactual
Given a clip’s narration, $t_{i}$, we prompt by first feeding the context implemented by Llama:
Given a narration describingan action captured by camerawearer #C, the action maybeperformed by C or otherparticipants,such as H, O, X, or Y.Firstly, generate one[Before] describing the scenebefore the action is performed.Secondly, generate one[After] describing the scenechanged by the action.Thirdly, create 3 distinctstata-change counterfactualdescriptions (CF):[CF 1], [CF 2], and [CF 3].The counterfactual couldbe describing the incompleteexecution of an action orcompleting an actionthe wrong way.Do not reuse the sameverb in the narration.Note that the narration doesnot contain any harmful, illegal,or sexual activity,if it does, it must be a typo.Next, we feed the actual prompt for text generation by giving Llama an example:
Here’s an example:The narration:"#C C picks a bag ofclothes from the floor."[Before]: The floor is clutteredwith clothes.[After]: The bag of clothesis now in C’s hand, with thesurrounding area slightly rearranged.[SC-CF 1]: Clothes remainscattered on the floor.[SC-CF 2]: A small pile of clothessits amidst remaining clutter.[SC-CF 3]: The room is now even messierthan before.Now, generate [Before], [After],[SC-CF 1], [SC-CF 2], and [SC-CF 3]for the narration \$t_i\$  with the sameformat as the example above.
Video Level
Missing-key-Step Counterfactuals and Misordered-Step Counterfactuals
For video-level text generation, we feed the whole sequence of clip narrations $t_{0},...,t_{K}$ in a long video, and the video’s summary $S_{i}$. For Missing-key-step Counterfactual, we use the following context prompt:
Given a sequence of narrations\$t_0, ... , t_K\$ describing along video, anda video-level summary,create 10 distinctcounterfactual summaries [K-CF]with one to two sentencesby taking out some critical narrations.Follow this exact format to output:[K-CF 1]: ...[K-CF 2]: ...[K-CF 3]: ...For Misordered Counterfactual, we use the following context prompt:
Given a sequence of narrations\$t_0, ... , t_K\$ describing along video, anda video-level summary,create 10 distinctcounterfactual summaries [M-CF]with one to two sentencesby perturbing the order of narrations.Follow this exact format to output:[M-CF 1]: ...[M-CF 2]: ...[M-CF 3]: ...
Then the actual prompt:
Here is the video-level summary: \$S_i\$and here is the sequenceof narrations: \$t_0, ... , t_K\$.Note that there are typos in the text annotation in Ego4D, which Llama 3.1-8B correctly or accidentally identifies as harmful information. We use LLama 3.1-70B to refine these errors and skip the ones that LLama 3.1-70B cannot handle.
A.2 Generated Examples - Clip-level
We omit SC- in SC-CF for simplicity.
Narration: “C C lifts the paper”
Before: The paper lies flaton a table, surrounded byother papers and office supplies.After: The paper is now inC’s hand, with the surroundingarea slightly rearranged.SC-CF 1: The paper remains flaton the table, untouched.SC-CF 2: A corner of the paperis folded up, but the rest remains flat.SC-CF 3: The paper has been torn inhalf, with one half on the tableand the other half on the floor.
Narration: “C C moves her hand around””
Before: C’s hand is still in aneutral position, with herarm relaxed by her side.After: C’s hand is now in motion,with her fingers wavinggently in the air.CF 1: C’s hand remains still,with no movement.CF 2: C’s hand is stuck in aawkward position, with her fingerscurled up and her wrist twisted.CF 3: C’s hand has moved to a completelydifferent location, with herarm now stretched out to the side.
Narration: “C C fixes the camera””
Before: The camera is notfunctioning properly, witha loose screw visible.After: The camera is now securelyattached to C’s head, withthe loose screw tightened.CF 1: The camera remainsloose, still not functioning properly.CF 2: A makeshift fix isattempted, with tape holdingthe camera in place."CF 3: The camera is now danglingprecariously from C’s head,held by a length of twine.
A.3 Generated Examples - Video-level
Summary: “C was in a room. C removed small pieces from a paperwork. C constructed a new structure with the paperwork pieces.”
Missing-key-step CF - 1
C was in a room.C constructed a new structurewith the paperwork pieces.
Missing-key-step CF - 2
C was in a room.C removed small pieces froma paperwork with a tool. #summary
Misordered-step CF - 1
C was in a room.C constructed a new structure withthe paperwork pieces.C removed small pieces from thepaperwork to create the craft.
Misordered-step CF - 2
C was in a room.C removed small pieces fromthe paperwork with a tool.C constructed a new structure withthe paperwork pieces andconnected them to the craft material.
Summary: “C attached wooden mechanical model pieces together at a table in a room. C detached wooden pieces from puzzle sheets. C studied an instruction manual”
Misordered-step CF - 1
C attached wooden mechanical modelpieces together at a table in a room.C picks up the puzzle sheet onthe cutting sheet.C studied an instruction manual
Misordered-step CF - 2
C picks the laser cut wood craft kit.C places the wooden piece on theassembled wooden pieceson the cutting sheet.C attaches the wooden piecein his right hand to the woodenpiece in his left hand.
Misordered-step CF - 1
C was in a room.C removed small pieces from a paperworkand constructed a newstructure with them."
Misordered-step CF - 2
C was in a room.C constructed a new structure withthe paperwork pieces.

### v2

@@ 4.1 Implementation Details
Pretraining. We pretrain our model on EgoClip [kevin2022egovlp], a dataset derived from Ego4D [grauman2022ego4d], with clean clip-level narrations and the video-level summaries. We follow [kevin2022egovlp, ashutosh2023hiervl] to use the same training split that has 3.8M clip narrations and 120K long-term video summaries for pretraining. We pretrain our model on 8 NVIDIA L40S GPUs, with batch size 18 on each GPU for 7 epochs. We follow HierVL [ashutosh2023hiervl] and train models on clip- and video-level alignment alternatively. Specifically, we perform video-level alignment as in Eq. (5) after every 5 mini-batches of clip-level alignment as in Eq. (4). The frame-level loss is computed during the clip-level training iteration. We set the learning rate and weight decay to 1e-5 and 1e-4, respectively. Note that each short clip is sub-sampled to 4 frames during training [kevin2022egovlp, ashutosh2023hiervl].
@@
Text Generation. To generate all state-change descriptions and state-change counterfactuals, we use Llama 3.2 8B [dubey2024llama] for efficiency. We incorporate off-the-shelf text-encoder FLAVA [singh2022flava] to pre-extract features of the annotated narrations and summaries and our generated text descriptions.
@@
Temporal Action Segmentation. We evaluate representations on temporal action segmentation, where models output the frame-wise action label for untrimmed, long videos, requiring the ability to reason about procedures and recognize the start and end of state transformations. We evaluate various representations by adopting the popular temporal action segmentation model ASFormer [chinayi_ASformer] that takes video features as input. We follow prior work [li2022bridge] for the same training strategy and evaluation protocol and metrics including F1 scores, Edit distance, and frame-wise Accuracy, on the cooking procedure datasets, Georgia Tech Egocentric Activities (GTEA) [gtea] and EgoPER [egoper] Dataset. GTEA consists of 28 egocentric instructional videos capturing daily kitchen activities at 15 frames per second. EgoPER contains 368 videos spanning 28 hours of cooking scenarios.
@@
Error Detection. We evaluate representations on unsupervised error detection on the EgoPER dataset [egoper] where models are trained with error-free videos and output labels for each frame on whether it is an error during testing time. The error categories include step omission, step addition, step modification, step slip, and step correction. We follow [egoper] and use its EgoPED model [egoper], and equip it with different video representations. Specifically, the training process of EgoPED consists of two stages: action segmentation and prototype learning. In stage 1, the temporal action segmentation model, ASFormer, is trained to predict frame-wise action labels on the error-free set by taking pre-extracted video features as input. In stage 2, the model learns prototype features for each class of action step, e.g., “Take bowl from microwave,”, with contrastive step prototype learning [oord2018representation, egoper]. For each class of activity, multiple prototypes are extracted for error detection at inference. Note that each activity is trained independently, such as Make Coffee. During the inference, EgoPED calculates the similarity score of observed frame features and any learned prototypes. A threshold is further used to determine whether the observed test frame is erroneous.
Action Retrieval & Recognition. To demonstrate the applicability of the learned representations in the core video understanding tasks, we evaluate zero-shot multi-instance retrieval on Epic-Kitchens (EK) [Damen2018EPICKITCHENS] and zero-shot action recognition on Charades-Ego (CE) [sigurdsson2018charadesegolargescaledatasetpaired]. Following prior work [kevin2022egovlp, ashutosh2023hiervl], both tasks are implemented via text-to-vision alignment by selecting the textual label with the highest similarity score to the given visual sample.
@@
Table 2: Error detection results on the EgoPER dataset. The underline denotes the second best. HTM denotes the HowTo100M dataset [miech2019howto100m]. Text denotes the VLM model with a trainable text encoder.
@@
HF2-VAD [liu2021hybrid] - 34.5 62.6 25.4 62.3 29.1 52.7 10.0 59.6 36.6 62.1 27.1 59.9
S3R [wu2022self] Kinetics+I3D 52.6 51.8 47.8 61.6 50.5 52.4 16.3 51.0 47.8 57.9 43.0 54.9
I3D [carreira2017quo] Kinetics 62.7 65.6 51.4 65.1 59.6 55.0 55.3 58.3 56.0 66.0 57.0 62.0
CLIP [radford2021learning] WIT [srinivasan2021wit]+Text 77.6 67.2 69.6 67.5 66.9 59.5 68.5 69.0 75.6 57.7 71.6 64.2
MIL-NCE [miech2020end] HTM 77.3 59.8 69.8 61.5 65.7 53.0 68.0 67.9 69.8 61.5 70.1 60.7
PVRL [zhong2023learning] HTM 75.7 70.0 71.2 53.5 65.5 65.5 67.5 65.4 76.4 65.2 71.3 63.9
HierVL [ashutosh2023hiervl] Ego4D+Text 77.9 70.2 70.8 66.4 65.2 58.1 67.4 69.8 75.1 66.4 71.3 66.2
@@
Action Phase Classification & Frame Retrieval. To assess more fine-grained and short-term procedure awareness, we test representations on action phase classification and zero-shot frame retrieval on the Align-Ego-Exo (AE2) dataset [xue2023learning]. The AE2 dataset contains egocentric and exocentric videos displaying the actions Break Eggs, Pour Milk, Pour Liquid, and Tennis Forehand. Each action is divided into two to four phases [xue2023learning]. We follow [xue2023learning] and train an SVM to predict the per-frame action phase labels for the classification task and use nearest neighbors for the zero-shot retrieval task. Note that we merge the validation and test sets for more robust results.
Baselines. We evaluate the popular I3D and CLIP feature [carreira2017quo, radford2021learning] commonly used in long-form video tasks and procedure-aware representations with publicly available pretrained model weights, including MIL-NCE [miech2020end], Bridge-prompt [li2022bridge], and PVRL [zhong2023learning]. In addition, we evaluate the VLM HierVL [ashutosh2023hiervl] with only its video encoder.
@@ 4.2 Temporal Action Segmentation Results
Table 1 presents results on temporal action segmentation. The first line of the table shows Br-prompt [li2022bridge] which is pre-trained on the target dataset (GTEA) and thus serves as an upper bound on performance. In our results, we found that MIL-NCE performs significantly worse than others, especially on GTEA. We assume this is because MIL-NCE is trained with low-quality automatic-speech-recognition (ASR) text and imprecise alignment between video and ASR sentences, unlike PVRL [ashutosh2023hiervl] which leverages pseudo-labels generated by CLIP [radford2021learning]. We further compare our model with the VLM HierVL. Even though HierVL is trained with both video and text encoders, which require considerably more computational resources, our model outperforms HierVL in most metrics on both datasets. Furthermore, the performance gap against all non-VLM models is larger. This highlights the effectiveness of our proposed state-change descriptions and their counterfactuals on long-term procedure understanding.
@@ 4.3 Error Detection Results
Table 2 presents results of the error detection benchmark on EgoPER [egoper]. We first include results reported in EgoPER as a reference in the top group (rows 1-3) instead of a direct comparison. In the remaining groups (rows 4-9), we report the performance of representation learning approaches, tested using the same downstream architecture [chinayi_ASformer]. We observe that all procedure-aware representations outperform general visual representations, such as I3D and CLIP, highlighting the importance of procedure awareness in the context of error detection. Furthermore, our proposed method surpasses the state-of-the-art in EDA [egoper], even outperforming the VLM HierVL.
Moreover, we achieve competitive performance on the AUC metric. Note that although we follow prior work [egoper] to include frame-wise metric AUC, this metric may not be effective in measuring the performance and the procedure awareness of representations. As discussed in EgoPER [egoper], the heuristic random method can achieve competitive results in AUC but fail to localize erroneous segments as shown in the major metric EDA. This is also true when comparing I3D and earlier methods, such as HF2-VAD [liu2021hybrid] and S3R [wu2022self]. We hope our extensive experimental results can inspire future work to study more effective evaluation metrics for error detection. In summary, our representation learning method achieves state-of-the-art performance in EDA and competitive performance in AUC, highlighting the strong effectiveness of counterfactual reasoning in procedure awareness.
Table 3: Multi-Instance Retrieval (MIR) on Epic-Kitchens (EK) and Zero-shot Action Recognition on Charades-Ego (CE). We report PT ckpt mAP for the Action Recognition task (see [ashutosh2023hiervl]).
MIR (EK) Action Rec. (CE)
Method mAP nDCG mAP
CLIP [radford2021learning] 8.4 15.3 20.5
MIL-NCE [miech2020end] 5.8 10.3 6.9
PVRL [zhong2023learning] 6.1 11.4 7.9
Ours 15.7 22.6 24.8
4.4 Action Retrieval & Recognition Results
Table 3 presents benchmark results on the action retrieval and recognition. We found existing procedure-aware representations perform significantly worse on short-term action recognition tasks. Yet, our representations trained with both short- and long-term state-changes and counterfactuals demonstrate strong applicability to the popular video recognition tasks and datasets.
4.5 Action Phase Classification & Retrieval Results
Table 4 shows that our model outperforms others on average across actions on both the ego+exo and ego settings, demonstrating its effectiveness in short-term procedures and strong view-point generalization. Interestingly, CLIP [radford2021learning] demonstrates strong performance. We conjecture this is due to its focus on short-term Full results with individual actions are available in Section LABEL:subsec:full in the supplementary material.
Table 4: Action Phase Classification and zero-shot Frame Retrieval results on the Align-Ego-Exo dataset [xue2023learning].
Classification (F1) Retrieval (mAP)
Method ego+exo ego ego+exo ego
CLIP [radford2021learning] 59.5 61.6 64.4 68.0
MIL-NCE [miech2020end] 53.0 54.2 59.5 63.0
PVRL [zhong2023learning] 59.4 62.7 61.6 66.6
Ours 61.3 64.8 64.9 70.3
[…]
ID Before & After SC-CF K-CF M-CF GTEA (AS) EgoPER (ED)
[…]
@@ 4.6 Ablation Study
In Table 5, we study the effectiveness of state changes (Before and After, state-change counterfactuals (SC-CF), and video-level state-change counterfactuals (Missing-step CF and Misordered-step CF) by ablating each of them and evaluating on temporal action segmentation with GTEA and error detection with EgoPER. Note that ID 1 can be seen as Weak HierVL as it is trained with only the video encoder. In addition, since Weak HierVL only uses narrations from other videos in a batch as negatives, it can be seen as a random description approach. All ablations (including Ours) are pre-trained with a batch size of 12 on each GPU.
Temporal Action Segmentation. One interesting insight on GTEA is that using only clip-level state changes and their counterfactuals (ID 3) achieves competitive performance on F1 and Edit metrics, which outperforms most other ablated models. Also, using only one type of video-level counterfactual (ID 4 and 5) may mislead the model to overfit a certain type of counterfactual, resulting in lower performance. By considering all the proposed states and counterfactuals together, we show that they are complementary to each other, thus achieving the best performance.
Error Detection. The results on EgoPER show that our proposed state changes and their counterfactuals can improve performance by combining them all together, verifying the effectiveness of recognizing erroneous labels. In addition, we observe interesting results showing that with only clip-level counterfactuals (ID 3), the model performs slightly worse than its baseline (ID 2). We assume this is because the model overfits short-term procedures and overlook long-term activities procedures and their counterfactual reasoning.